Let's assume the following data represents the time from when a patient is discharged from North Shore Hospital to the time the discharged patient's bed is ready to be assigned to a new patient is referred to as the bed assignment turnaround time. If the bed turnaround time is excessive, it can cause problems with patient flow and delay medical procedures throughout the hospital. This can cause long waiting times for physicians and patients, thus creating customer dissatisfaction. The admissions RN has assigned a patient care associate to measure the bed turnaround time for a randomly selected bed each morning, afternoon, and evening for 30 days. Following are the bed turnaround time sample observations.

QUESTION: Comment on whether or not a requirement that bed turnaround time fall within 80 - 200 minutes is feasible. Support your answer using Cp as a measure of capability

167, 127, 135, 128, 97,

112, 140, 155, 122, 223,

167, 145, 84, 146, 169,

216, 190, 178, 103, 115

104, 122, 105, 167, 98

125, 135, 115, 105, 140

SHOW WORK
Answer #1

Given : n=30 , USL=200 and LSL=80

We calculate the sample standard deviation as ,

sample standard deviation=STDEV(D2:H7)=34.70

Process standard deviation=Sample standard deviation/sqrt(n)=E11/SQRT(30)=6.335

Process capability ratio =Cp =(USL-LSL)/6*process standard deviation

=(200-80)/6*6.335

=(200-80)/(6*E13)

=3.157

A value of Cp >= 1.33 indicates that the process is capable .Hence,the requirement that bed turnaround time fall within 80-200 is feasible.